### v1

@@ 1.1 Basic Definitions and Our Contributions
Before we formally state our main result, we first introduce some definitions. Given a graph $G=(V,E,\mathbf{w})$ with weight vector $\mathbf{w}\in\mathbb{R}_{+}^{\binom{V}{2}}$, and a graph $M=(V_{M},E_{M})$ which we assume to be a frequently occurring subgraph of $G$ and which is referred to as a motif. An instance of motif $M$ is a subgraph of $G$, which is isomorphic to $M$, and the weight of an instance $I=(V_{I},E_{I})$ is defined222 As justified in KMS+ (22), in integer-weighted graphs, motifs are often viewed as unweighted multigraphs, where each edge is replaced by multiple copies based on its weight. Extending this idea, the motif weight is naturally defined as the product of its edge weights. as the product of its edge weights, i.e., $w(I)=\prod_{e\in E_{I}}w(e)$.
@@
We consider the standard notion of edge privacy and two graphs are called neighboring if the two vectors corresponding to their edge weights differ by at most $1$ in the $\ell_{1}$ norm (see Definition 5 for the formal definition).
@@
Note that for unweighted graphs with $m$ edges, it holds that $W=m$ and $w_{\max}=1$. Thus, our algorithm achieves $\tilde{O}(\sqrt{m\cdot\ell_{3}(G)}\cdot n/\varepsilon^{3/2})$ additive error. Note that for certain classes of graphs $G$, $\ell_{3}(G)$ can be relatively small (in comparison to $n$). For example, if $G$ has maximum degree at most $d$, then $\ell_{3}(G)\leq d$. (See Section 5 for more discussions of local sensitivity.) Our bound can be compared to the trivial upper bounds $O(m^{3/2}),O(n^{3})$ or $O(mn)$ on the triangle-motif size of any cut in a graph. We remark that before our work, there was no known polynomial-time DP algorithm for this problem. We also give a $(\varepsilon,0)$-DP algorithm that outputs a synthetic graph (in Appendix C) with error $\tilde{O}(n^{\frac{5}{2}})$ based on randomized response, using an analysis similar to GRU (12) for the edge case. Note that the bound $O(\sqrt{m\cdot\ell_{3}(G)}\cdot n)$ from Theorem 3 is never worse than, and can often be much better than, the $\tilde{O}(n^{\frac{5}{2}})$ bound from randomized response.
Our algorithms are built upon the framework of solving a related convex problem using the mirror descent approach and adding noise appropriately, similar to the approach in EKKL (20). However, unlike the edge case, answering motif cut queries on a graph does not align with the extensively studied problem of query release for exponentially sized families of linear queries on a dataset, due to the inherently non-linear nature of motif cut queries. Consequently, we had to resolve several technical challenges (see Section 1.3).
@@
We remark that by the proof of the above lower bound, we can also obtain a lower bound of $\Omega(\sqrt{\frac{mn}{\varepsilon}}\ell_{3}(G)\cdot(1-c))$ for unweighted graphs which may contain multiple edges (see the proof of Theorem 44). Furthermore, our lower bound also generalizes to all $K_{h}$ motifs, i.e., the complete graph on $h$ vertices, for any constant $h\geq 2$. For the $K_{h}$ motifs, our lower bound on the additive error becomes $\Omega\left(\max\left({\frac{\sqrt{mn}\cdot\ell_{h}(G)}{\varepsilon}}(1-c),{\frac{\sqrt{Wn}\cdot\ell_{h}(G)}{\sqrt{\varepsilon}}}(1-c)\right)\right)$ (while the corresponding requirement for $p$ is $((\log n)/n)^{1/(h-1)}\ll p\leq\frac{1}{2}$), where $\ell_{h}(G)$ is the local sensitivity of $K_{h}$-motif cuts of $G$ (see Theorem 44). Note that it also recovers the $\Omega(\sqrt{mn})$ lower bound for the edge motif case (as here $\ell_{2}(G)=1$). Our lower bound is based on the connections to discrepancy theory shown by MN (12) and a lower bound on the discrepancy of some coloring function on hypergraphs.
@@ 1.3 Our Techniques
Directly applying the EKKL approach to the associated hypergraph or the triangle-motif weighted graph do not seem work due to the non-linearity property of triangle cut. That is, the sum of the triangle-motif sizes for cut $(S,V\setminus S)$ of two graphs $G_{1}$ and $G_{2}$, is not equal to the triangle-motif size for cut $(S,V\setminus S)$ of the sum of the two graphs $G_{1}+G_{2}$, or formally, $\mathrm{Cut}_{\triangle}^{(G_{1})}(S,V\setminus S)+\mathrm{Cut}_{\triangle}^{(G_{2})}(S,V\setminus S)\neq\mathrm{Cut}_{\triangle}^{(G_{1}+G_{2})}(S,V\setminus S)$. We compare in more details about the difference of the EKKL approach and ours in Section 1.3 and also discuss some tempting but unsuccessful approaches in Section A.2.
@@
where $\overline{\mathbf{A}}_{\triangle}$ is the triangle adjacency matrix of the input graph $\bar{G}=(V,E,\overline{\mathbf{w}})$ with sum of edge weights at most $W$. Intuitively, finding a weight vector $\mathbf{w}$ that minimizes the inner SDP of Equation 1 will result in a synthetic graph with a low triangle-motif cut difference between $\mathbf{w}$ and $\overline{\mathbf{w}}$. It is crucial that the objective function is defined in terms of the weight function $\mathbf{w}$ of the target graph $G$ instead of $\mathbf{w}_{\triangle}$, which is the edge weight vector of the triangle-motif weighted graph ${G}_{\triangle}$.
However, due to the non-linearity property of triangle-motif cut, Equation 1 is no longer convex with respect to $\mathbf{w}$, thus it cannot be solved by convex optimization techniques. Our approach is to add a convexity regularizer [^4] Here, we use the term $6nw_{\max}\sum_{e\in\binom{V}{2}}(\mathbf{w}_{e}-\overline{\mathbf{w}}_{e})^{2}$ to provide a basic understanding of the main idea; in our actual proof, we incorporate a slightly more complicated term to achieve a smaller error. $6nw_{\max}\sum_{e\in\binom{V}{2}}(\mathbf{w}_{e}-\overline{\mathbf{w}}_{e})^{2}$ to the objective function to control the convexity, while ensuring that we do not add too much error. Define
@@
With the convexity regularizer, we can apply Danskin’s theorem Dan (12) to show that our objective function $f_{\triangle}(\mathbf{w}):=\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X})$ is convex. However, solving the corresponding optimization using the stochastic mirror descent becomes more challenging. Specifically, we need to utilize the gradient of $f_{\triangle}(\mathbf{w})$ for updating the mirror descent, which corresponds to finding the minimum of the Bregman divergence associated to some mirror map function (see Section 2.4), and bounding the final additive error. Unlike the edge case, where the gradient is a constant, the gradient of $f_{\triangle}(\mathbf{w})$ is a relatively complex function that depends on $\mathbf{w}$ (see Lemma 26). This dependency arises from the higher-order nature of the triangle-motif and introduces additional technical difficulties.
Our solution is to impose additional constraints $\{\mathbf{w}_{e}\leq w_{\max},e\in\binom{V}{2}\}$ on $\mathcal{X}^{\prime}$ to obtain a more restricted domain $\mathcal{X}$, and then introduce a new mirror descent update rule based on a greedy method. Specifically, we transform the problem of updating $\mathbf{w}$, i.e., finding the minimizer of the Bregman divergence, into another convex optimization problem (see Equation 19). Utilizing the Karush-Kuhn-Tucker (KKT) conditions GGKC (21), we derive the necessary and sufficient conditions that the optimal solution of this optimization problem must satisfy. Based on these conditions, we develop an efficient algorithm (Algorithm 3) to update the weight vector $\mathbf{w}$ using a greedy method.
@@
Our lower bound is established within the discrepancy framework introduced by Muthukrishnan and Nikolov MN (12). This framework was also utilized in proving the lower bound for DP algorithms concerning edge motif cut in EKKL (20). Essentially, if there exists a DP algorithm $\mathcal{M}$ for the motif cut problem with additive error smaller than the discrepancy of the motif cut size function over certain classes of graphs, then algorithm $\mathcal{M}$ can be exploited to approximately recover the input, thereby compromising privacy and resulting in a contradiction. To leverage this framework effectively, we need to show that the discrepancy of this function over some class of graphs is relatively large to exclude DP algorithms with small additive error. Specifically, to derive the lower bound for the $K_{h}$-motif cut for any complete graph $K_{h}$ with $h$ vertices, we employ the discrepancy of $3$-colorings (i.e. each motif is colored with $+1,-1$ or 00) of $h$-uniform hypergraphs, which is a generalization of the corresponding discrepancy of graphs. In particular, we identify a set of properties (e.g., each vertex has roughly the same degree and each edge belongs to roughly the same number of $K_{h}$ instances) sufficient to ensure that: 1) a graph from $G(n,p)$ will satisfy with high probability, and 2) the corresponding $h$-uniform hypergraph (defined by treating any subset of $h$ vertices as a hyperedge) exhibits large discrepancy.
@@ 2.2 Motif Cut and Motif Adjacency Matrix
In this paper, we focus on the special case when the motif is triangle. The “$M$” in the relevant notation will be replaced by $\triangle$”.
@@
Let $\mathbf{E}_{\triangle}^{((i,j),(k,\ell))}$ denote the second-order derivative of $\mathbf{A}_{\triangle}$ at $\mathbf{w}_{(i,j)}$ and $\mathbf{w}_{(k,\ell)}$, and let $\mathbf{B}_{(e)}$ denote a matrix which has value $1$ at the entry corresponding to $e$ and value 00 otherwise. Then it holds that
$\left(\mathbf{D}_{\triangle}^{(k,\ell)}\right)_{i,j}=\begin{cases}\sum_{s\neq k,\ell}\mathbf{w}_{(k,s)}\mathbf{w}_{(s,\ell)},&(k,\ell)=(i,j)\\ \mathbf{w}_{(k,j)}\mathbf{w}_{(j,\ell)},&k=i\ and\ j\neq\ell\\ 0,&o.w.\\ \end{cases}$
$\mathbf{E}_{\triangle}^{((i,j),(k,\ell))}=\begin{cases}\mathbf{w}_{(j,\ell)}\cdot(\mathbf{B}_{(i,j)}+\mathbf{B}_{(i,\ell)}+\mathbf{B}_{(j,\ell)}),&i=k\ and\ j\neq\ell\\ 0,&o.w.\\ \end{cases}$
@@ 2.4 Convex Optimization
Output: $\hat{x}\in\mathcal{X}$ s.t. $f(\hat{x})$ approximates $\min_{x\in\mathcal{X}}f(x)$.
@@
Denote $R_{\ell,t}$ as the randomness while computing $g^{(t)}$ in the outer iteration $\ell$. For any fixed outer iteration $\ell$, the difference between $\mathrm{E}_{\{\gamma_{\ell},R_{\ell,t}\}}[f(\hat{x}_{\ell})]$ and $\min_{x\in\mathcal{X}}f(x)$ is bounded by the following theorem. Here, $\mathrm{E}_{\{\gamma_{\ell},R_{\ell,t}\}}$ represents the expectation taken over the randomness corresponding to random variables $\gamma_{\ell},R_{\ell,t}$, where $\ell=1,\dots,L$ and $t=1,\dots,T$.
@@
[…]
@@
[…]
@@
$\mathrm{s.t.}$ $g_{i}(x)\leq 0,\forall i\in\{1,\dots,m_{1}\}$

### v2

@@ 1.1 Basic Definitions and Our Contributions
Before we formally state our main result, we first introduce some definitions. Given a graph $G=(V,E,\mathbf{w})$ with weight vector $\mathbf{w}\in\mathbb{R}_{+}^{\binom{V}{2}}$, and a graph $M=(V_{M},E_{M})$ which we assume to be a frequently occurring subgraph of $G$ and which is referred to as a motif. An instance of motif $M$ is a subgraph of $G$, which is isomorphic to $M$, and the weight of an instance $I=(V_{I},E_{I})$ is defined [^2] As justified in KMS+ (22), in integer-weighted graphs, motifs are often viewed as unweighted multigraphs, where each edge is replaced by multiple copies based on its weight. Extending this idea, the motif weight is naturally defined as the product of its edge weights. as the product of its edge weights, i.e., $w(I)=\prod_{e\in E_{I}}w(e)$.
@@
We consider the standard notion of edge privacy and two graphs are called neighboring if the two vectors corresponding to their edge weights differ by at most 11 in the $\ell_{1}$ norm (see Definition˜5 for the formal definition).
@@
Note that for unweighted graphs with $m$ edges, it holds that $W=m$ and $w_{\max}=1$. Thus, our algorithm achieves $\tilde{O}(\sqrt{m\cdot\ell_{3}(G)}\cdot n/\varepsilon^{3/2})$ additive error. Note that for certain classes of graphs $G$, $\ell_{3}(G)$ can be relatively small (in comparison to $n$). For example, if $G$ has maximum degree at most $d$, then $\ell_{3}(G)\leq d$. (See Section˜5 for more discussions of local sensitivity.) Our bound can be compared to the trivial upper bounds $O(m^{3/2}),O(n^{3})$ or $O(mn)$ on the triangle-motif size of any cut in a graph. We remark that before our work, there was no known polynomial-time DP algorithm for this problem. We also give a $(\varepsilon,0)$-DP algorithm that outputs a synthetic graph (in Appendix˜C) with error $\tilde{O}(n^{\frac{5}{2}})$ based on randomized response, using an analysis similar to GRU (12) for the edge case. Note that the bound $O(\sqrt{m\cdot\ell_{3}(G)}\cdot n)$ from Theorem˜3 is never worse than, and can often be much better than, the $\tilde{O}(n^{\frac{5}{2}})$ bound from randomized response.
Our algorithms are built upon the framework of solving a related convex problem using the mirror descent approach and adding noise appropriately, similar to the approach in EKKL (20). However, unlike the edge case, answering motif cut queries on a graph does not align with the extensively studied problem of query release for exponentially sized families of linear queries on a dataset, due to the inherently non-linear nature of motif cut queries. Consequently, we had to resolve several technical challenges (see Section˜1.3).
@@
We remark that by the proof of the above lower bound, we can also obtain a lower bound of $\Omega(\sqrt{\frac{mn}{\varepsilon}}\ell_{3}(G)\cdot(1-c))$ for unweighted graphs which may contain multiple edges (see the proof of Theorem˜44). Furthermore, our lower bound also generalizes to all $K_{h}$ motifs, i.e., the complete graph on $h$ vertices, for any constant $h\geq 2$. For the $K_{h}$ motifs, our lower bound on the additive error becomes $\Omega\left(\max\left({\frac{\sqrt{mn}\cdot\ell_{h}(G)}{\varepsilon}}(1-c),{\frac{\sqrt{Wn}\cdot\ell_{h}(G)}{\sqrt{\varepsilon}}}(1-c)\right)\right)$ (while the corresponding requirement for $p$ is $((\log n)/n)^{1/(h-1)}\ll p\leq\frac{1}{2}$), where $\ell_{h}(G)$ is the local sensitivity of $K_{h}$-motif cuts of $G$ (see Theorem˜44). Note that it also recovers the $\Omega(\sqrt{mn})$ lower bound for the edge motif case (as here $\ell_{2}(G)=1$). Our lower bound is based on the connections to discrepancy theory shown by MN (12) and a lower bound on the discrepancy of some coloring function on hypergraphs.
@@ 1.3 Our Techniques
Directly applying the EKKL approach to the associated hypergraph or the triangle-motif weighted graph do not seem work due to the non-linearity property of triangle cut. That is, the sum of the triangle-motif sizes for cut $(S,V\setminus S)$ of two graphs $G_{1}$ and $G_{2}$, is not equal to the triangle-motif size for cut $(S,V\setminus S)$ of the sum of the two graphs $G_{1}+G_{2}$, or formally, $\mathrm{Cut}_{\triangle}^{(G_{1})}(S,V\setminus S)+\mathrm{Cut}_{\triangle}^{(G_{2})}(S,V\setminus S)\neq\mathrm{Cut}_{\triangle}^{(G_{1}+G_{2})}(S,V\setminus S)$. We compare in more details about the difference of the EKKL approach and ours in Section˜1.3 and also discuss some tempting but unsuccessful approaches in Section˜A.2.
@@
where $\overline{\mathbf{A}}_{\triangle}$ is the triangle adjacency matrix of the input graph $\bar{G}=(V,E,\overline{\mathbf{w}})$ with sum of edge weights at most $W$. Intuitively, finding a weight vector $\mathbf{w}$ that minimizes the inner SDP of Equation˜1 will result in a synthetic graph with a low triangle-motif cut difference between $\mathbf{w}$ and $\overline{\mathbf{w}}$. It is crucial that the objective function is defined in terms of the weight function $\mathbf{w}$ of the target graph $G$ instead of $\mathbf{w}_{\triangle}$, which is the edge weight vector of the triangle-motif weighted graph ${G}_{\triangle}$.
However, due to the non-linearity property of triangle-motif cut, Equation˜1 is no longer convex with respect to $\mathbf{w}$, thus it cannot be solved by convex optimization techniques. Our approach is to add a convexity regularizer [^4] Here, we use the term $6nw_{\max}\sum_{e\in\binom{V}{2}}(\mathbf{w}_{e}-\overline{\mathbf{w}}_{e})^{2}$ to provide a basic understanding of the main idea; in our actual proof, we incorporate a slightly more complicated term to achieve a smaller error. $6nw_{\max}\sum_{e\in\binom{V}{2}}(\mathbf{w}_{e}-\overline{\mathbf{w}}_{e})^{2}$ to the objective function to control the convexity, while ensuring that we do not add too much error. Define
@@
With the convexity regularizer, we can apply Danskin’s theorem Dan (12) to show that our objective function $f_{\triangle}(\mathbf{w}):=\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X})$ is convex. However, solving the corresponding optimization using the stochastic mirror descent becomes more challenging. Specifically, we need to utilize the gradient of $f_{\triangle}(\mathbf{w})$ for updating the mirror descent, which corresponds to finding the minimum of the Bregman divergence associated to some mirror map function (see Section˜2.4), and bounding the final additive error. Unlike the edge case, where the gradient is a constant, the gradient of $f_{\triangle}(\mathbf{w})$ is a relatively complex function that depends on $\mathbf{w}$ (see Lemma 26). This dependency arises from the higher-order nature of the triangle-motif and introduces additional technical difficulties.
Our solution is to impose additional constraints $\{\mathbf{w}_{e}\leq w_{\max},e\in\binom{V}{2}\}$ on $\mathcal{X}^{\prime}$ to obtain a more restricted domain $\mathcal{X}$, and then introduce a new mirror descent update rule based on a greedy method. Specifically, we transform the problem of updating $\mathbf{w}$, i.e., finding the minimizer of the Bregman divergence, into another convex optimization problem (see Equation˜11). Utilizing the Karush-Kuhn-Tucker (KKT) conditions GGKC (21), we derive the necessary and sufficient conditions that the optimal solution of this optimization problem must satisfy. Based on these conditions, we develop an efficient algorithm (Algorithm˜3) to update the weight vector $\mathbf{w}$ using a greedy method.
@@
Our lower bound is established within the discrepancy framework introduced by Muthukrishnan and Nikolov MN (12). This framework was also utilized in proving the lower bound for DP algorithms concerning edge motif cut in EKKL (20). Essentially, if there exists a DP algorithm $\mathcal{M}$ for the motif cut problem with additive error smaller than the discrepancy of the motif cut size function over certain classes of graphs, then algorithm $\mathcal{M}$ can be exploited to approximately recover the input, thereby compromising privacy and resulting in a contradiction. To leverage this framework effectively, we need to show that the discrepancy of this function over some class of graphs is relatively large to exclude DP algorithms with small additive error. Specifically, to derive the lower bound for the $K_{h}$-motif cut for any complete graph $K_{h}$ with $h$ vertices, we employ the discrepancy of $3$-colorings (i.e. each motif is colored with +1,1+1,-1 or 0) of $h$-uniform hypergraphs, which is a generalization of the corresponding discrepancy of graphs. In particular, we identify a set of properties (e.g., each vertex has roughly the same degree and each edge belongs to roughly the same number of $K_{h}$ instances) sufficient to ensure that: 1) a graph from $G(n,p)$ will satisfy with high probability, and 2) the corresponding $h$-uniform hypergraph (defined by treating any subset of $h$ vertices as a hyperedge) exhibits large discrepancy.
@@ 2.2 Motif Cut and Motif Adjacency Matrix
In this paper, we focus on the special case when the motif is triangle. The “MM” in the relevant notation will be replaced by $\triangle$”.
@@
Let $\mathbf{E}_{\triangle}^{((i,j),(k,\ell))}$ denote the second-order derivative of $\mathbf{A}_{\triangle}$ at $\mathbf{w}_{(i,j)}$ and $\mathbf{w}_{(k,\ell)}$, and let $\mathbf{B}_{(e)}$ denote a matrix which has value $1$ at the entry corresponding to $e$ and value 0 otherwise. Then it holds that
[…]
@@ 2.4 Convex Optimization
Output: $\hat{x}\in\mathcal{X}$ s.t. f(x^)f($\hat${x}) approximates $\min_{x\in\mathcal{X}}f(x)$.
@@
[…]
@@
Let $\Phi$ be a $\rho$-strongly convex map with respect to $\|\cdot\|$. Given a convex function $f$ defined over convex set $\mathcal{X}$ with $x^{*}=\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x)$. Additionally, assume the noises satisfy $\mathbb{E}\left[\gamma_{\ell}\right]=0$. Consider Algorithm˜1 for some fixed outer iteration $\ell\leq L$.
Assume that $\Phi(x^{*})-\min_{x\in\mathcal{X}}\Phi(x)=R^{2}$, $\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}[g^{(t)}]=\nabla f(x^{(t)})$ and $\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}[\|g^{(t)}\|_{*}^{2}]\leq B^{2}$ for all $t\leq T$, where $\|\cdot\|_{*}$ denotes the norm dual to $\|\cdot\|$. After $T$ iterations with step length $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$, denote the output of the Line 3 to Line 11 in Algorithm˜1 as $\hat{x}_{\ell}\in\mathcal{X}$, then $\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}[f(\hat{x}_{\ell})]\leq f(x^{*})+RB\sqrt{\frac{2}{\rho T}}$.
@@
$\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}\left[f(\frac{1}{T}\sum_{t=1}^{T}x^{(t)})-f(x)\right]\leq\frac{1}{T}\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}\left[\sum_{t=1}^{T}(f(x^{(t)})-f(x))\right]$
$\leq$ $\frac{1}{T}\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}\left[\sum_{t=1}^{T}\nabla f(x^{(t)})^{\top}(x^{(t)}-x)\right]=\frac{1}{T}\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}\left[\sum_{t=1}^{T}\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}\left[g^{(t)}\right]^{\top}(x^{(t)}-x)\right]$
$=$ $\frac{1}{T}\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}\left[\sum_{t=1}^{T}{g^{(t)}}^{\top}(x^{(t)}-x)\right]\leq\frac{1}{T}\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}\left[\frac{R^{2}}{\eta}+\frac{\eta}{2\rho}\sum_{t=1}^{T}\left\|g^{(t)}\right\|_{*}^{2}\right]\leq\frac{R^{2}}{\eta}+\frac{\eta B^{2}}{2\rho}.$
@@
Let $\Phi$ be a $\rho$-strongly convex map with respect to $\|\cdot\|$. Given a convex function $f$ defined over convex set $\mathcal{X}$ with $x^{*}=\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x)$. Assume that $\Gamma_{\ell}$ has zero expectation for each $\ell$. Assume that $\Phi(x^{*})-\min_{x\in\mathcal{X}}\Phi(x)=R^{2}$, $\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}[g^{(t)}]=\nabla f(x^{(t)})$ and $\mathbb{E}_{\{\gamma_{\ell},R_{\ell,t}\}}[\|g^{(t)}\|_{*}^{2}]\leq B^{2}$ for all $t\leq T$ and $\ell\leq L$, where $\|\cdot\|_{*}$ denotes the norm dual to $\|\cdot\|$. Algorithm˜1 with parameter $\eta=\frac{R}{B}\sqrt{\frac{2}{T}}$ will output $\hat{x}\in\mathcal{X}$ s.t. $f(\hat{x})\leq 3f(x^{*})+3RB\sqrt{\frac{2}{\rho T}}$ with probability at least $1-\beta$.
@@
$\mathrm{s.t.}\ \$$g_{i}(x)\leq 0,\forall i\in\{1,\dots,m_{1}\}$